4. Find the singular value decomposition of matrix 3 6 -3 -6 B 23 -4 17 -16
Answer #1

-3 2 3 =G- -4 -16-3 - 6 23-4 -3 - 6 2 3A B: [836 -32g 344 - 328 2 8 363 44 \1 80 A1 72 = 180,0 0 0 (120-2) = 120 + 18 Q 0 0 g